2) Assume that the price elasticity of demand is 1.5 and supply is perfectly inelastic. And



suppose that the supply of the Palestinian olive oil has increased by 30%. What is the new



price if the old price is $10/KG? Explain.

1
Expert's answer
2021-12-05T18:48:21-0500

Solution:

Price elasticity of demand (PED) = "\frac{\%\;change\; in\; quantity\; demanded}{\%\; change\; in\; price}"

PED = 1.5


"\%\;change\; in\; quantity\; demanded = 30\%"

"Let\; \%\;change\; in\; price\;be= P"

PED = 1.5


1.5 = "\frac{30\%}{P}"


P = "\frac{30\%}{1.5} = 20\%"


Price has decreased by 20%

Previous price = $10

Price reduction by "20\%" = ("100\% - 20\%) \times10 = \$8"

The new price is "\$8"